OpenStudy (anonymous):

Use sum or difference formula to find: cos 5pi/7cos8pi/21 + sin 5pi/7 sin8pi/21

OpenStudy (jdoe0001):

\(\bf cos(\alpha-\beta)=cos(\alpha)cos(\beta)+sin(\alpha)sin(\beta)\) http://d2gbom735ivs5c.cloudfront.net/m/geometry/images/323d3f359a7b0b76ce0d94ea3b90ffd6.png